### v2

# Matching with Generalized Sequential Reserves

Abstract.
This paper introduces a general theory of sequential reserves by formulating a parameterized family of generalized sequential (GS) choice rules for institutions consisting of multiple divisions (or reserve categories). We show that the cumulative offer mechanism (COM) is the unique stable and strategy-proof mechanism when institutions’ choice rules are in the GS family. Choice rules in real-life market design applications appear naturally in the GS family. We present two such real-world applications: affirmative action implementations in India and Chinese high school admissions.
This paper subsumes and replaces the earlier working paper titled “Matching with Generalized Lexicographic Choice Rules.” We are grateful to the editor, Faruk Gul, the associate editor, and the two reviewers for their constructive and helpful feedback. We also thank Itai Ashlagi, Eduardo Azevedo, David Delacrétaz, Battal Doğan, Federico Echenique, Aytek Erdil, Isa Hafalır, Onur Kesten, Aditya Kuvalekar, Kriti Manocha, Debasis Mishra, Josué Ortega, Assaf Romm, Al Roth, Arunava Sen, Bumin Yenmez, Kemal Yildiz, as well as the participants at the ISU Market Design Conference, the 2022 Asian Meeting of the Econometric Society in East and Southeast Asia, the 2022 Conference on Mechanism and Institution Design, the 2024 North American Meeting of the Econometric Society, and the 2024 Asian Meeting of the Econometric Society for their helpful comments and suggestions.
1. Introduction
Many real-life allocation problems involve matching individuals with institutions that are made up of multiple divisions such as firms, hospitals, or schools. In rationing healthcare resources (Pathak, Sönmez, Ünver, and Yenmez (2024) and Aziz and Brandl (2024)) and the allocation of public goods such as seats in prestigious engineering schools (Baswana, Chakrabarti, Chandran, Kanoria, and Patange (2019), Aygün and Turhan (2022)) and government jobs (Sönmez and Yenmez (2022)) available resources are divided into different reserve categories, each with a subset of individuals eligible for that category. Divisions (or reserve categories) are endowed with their own choice rules. In some applications, divisions’ choice rules are induced from a strict ranking of alternatives, by always selecting the q-best elements whenever available. Such choice rules are called q-responsive. Similarly, in many rationing practices, there is a baseline priority ordering on applicants, and resource units are assigned according to this baseline priority order in each reserve category, also depending on the eligibility of candidates. However, in many other allocation or rationing practices, divisions’ choices are more complicated. To make it precise, consider a division that selects applicants on the basis of their merit scores but is also required to select minimum numbers of disabled individuals and women applicants, whenever possible. This choice rule is obviously not responsive.
Often, divisions are interconnected in that the number of positions available in one division depends on how many applicants other divisions hire. This creates inter-divisional constraints across the institution. For example, consider a business school whose departments, such as marketing, economics, and finance, are in the job market to hire new faculty. Suppose that the marketing department hires first, followed by economics, which is then followed by finance. The number of available positions for finance might be increased if either the marketing or the economics department hires fewer faculty than they originally planned, given that the total budget of the school is not exceeded. Roth and Peranson (1999) were the first to address interdivisional constraints of this kind in matching theory. They introduced the notion of “reversion" in the context of the National Residency Matching Program (NRMP), which matches doctors with hospital programs. Reversion is a practice where unfilled positions in one hospital program can be transferred to other programs while respecting the hospital’s overall capacity limits. In their analysis, Roth and Peranson (1999) identified reversion as a significant complication in the NRMP’s matching process.
There exists a significant body of literature that examines the design and implementation of soft reserve policies in rationing and resource allocation problems. The specific method of soft reservation implementation has important distributional consequences. In the context of admissions to technical universities in India, Aygün and Turhan (2022) introduced a soft reserve policy that reverts unfilled slots reserved for particular groups to the open category, which is filled before reserve categories by updating the position categories’ capacities and re-running the choice rule from scratch. In a similar framework, Aygün and Turhan (2025) developed forward transfer choice rules in which reserved slots that are not filled are allocated as open category after the reserve categories are filled. Hu and Zhang (2025) present various soft reserve policies specifically currently in use for high school admissions in China.
This paper studies the choice problem of multi-divisional institutions that face cross-divisional constraints. We introduce a family of choice rules termed generalized sequential (GS) rules, which incorporate two fundamental generalizations. First, these rules allow divisions to employ substitutable and size-monotonic choice rules rather than constraining them to simpler choice rules based on given ranking of alternatives. Second, these rules facilitate a more comprehensive implementation of soft reserves, as defined through transfer functions. Specifically, we present a parameterized family of choice rules constructed from choice rules that satisfy the substitutes property and size monotonicity—two of the most prominent properties in the market design literature. Although choice rules in the GS family may fail the substitute condition and size monotonicity, we show in our Theorem 1 that every choice rule in the GS family has a completion that satisfies the substitute condition and size monotonicity. A direct corollary of Theorem 1 is that the generalized deferred acceptance mechanism is the unique mechanism that is stable and strategy-proof in centralized markets.
The choice rules in the GS family have the following structure: Each institution has a number of divisions, each of which has a certain number of available positions. There is a predetermined order of precedence over divisions according to which the institution fills the positions of its divisions. Each division is endowed with its own choice rule. A division’s choice rule has two inputs: the set of available options and its capacity. Both inputs depend on the choices made by the preceding divisions. The set of available options can be thought of as the set of remaining options from the choices of divisions that precede it. The capacity of each division is an increasing function of the number of unfilled positions of the divisions that precede it, given by an exogenously specified transfer function. The aggregate choice rule of an institution is then defined as the union of its divisions’ choice rules. The collection of divisions’ choice rules and the transfer function identify an aggregate choice rule.
The choice rule for each division satisfies the substitutes and the size monotonicity properties. The substitutes condition (Kelso and Crawford (1982) and Gul and Stacchetti (1999)) and size monotonicity (Alkan and Gale (2003) and Hatfield and Milgrom (2005)) impose restrictions on the choice function when the set of available alternatives changes while the capacity is fixed. A choice rule satisfies the substitute property if there are no two alternatives $x$ and $y$ that are complementary, in the sense that gaining access to $x$ makes $y$ desirable. A choice rule is size monotonic if it chooses weakly more alternatives whenever the set of available alternatives expands.
The choice rule of each division also satisfies the quota monotonicity property, which is similar to the notion of choice rule expansion defined in Chambers and Yenmez (2017). [^1] The quota monotonicity property is a stronger condition than the expansion property of Chambers and Yenmez (2017). The quota monotonicity property imposes restrictions on the choice rule when the capacity of the division changes, while the set of available alternatives remains unchanged. A choice rule satisfies quota monotonicity if
  • all alternatives selected under the initial capacity are chosen when the capacity increases, and
  • the difference between the number of chosen alternatives under the increased and initial capacities cannot exceed the increase in capacity.
Divisions’ choice rules with initial and increased capacities can be regarded as two different choice rules that satisfy the substitutes condition and size monotonicity. The quota monotonicity property imposes a tight relationship between them. The choice rule with the increased capacity is an expansion of the choice rule with the initial capacity.
We formulate cross-division constraints via transfer functions. Given the order of precedence over divisions, the capacity of a division is an increasing function of the number of vacancies in the divisions that precede it. We require that the transfer function be monotone (Westkamp (2013)). That is, it satisfies the following conditions: $(1)$ whenever there are weakly more seats unfilled in every division preceding the $j^{th}$ division, weakly more slots should be available for the $j^{th}$ division, and $(2)$ an institution cannot excessively decrease the capacity of division $j$ in response to increased demand in some divisions processed earlier.
An aggregate choice rule is in the GS family if divisions’ choice rules satisfy the substitutes condition, size monotonicity, and quota monotonicity, and the transfer function is monotonic. Aggregate choice rules in the GS family may fail to satisfy the substitutes property and size monotonicity. In our Theorem 1, we show that each aggregate choice rule in the GS family has a completion that satisfies the substitutes property and size monotonicity (Hatfield and Kominers (2019)). An immediate corollary of Theorem 1 is that the generalized deferred acceptance mechanism (or cumulative offer mechanism) is the unique mechanism that is stable and strategy-proof.
In practice, policymakers frequently implement choice rules characterized by intuitive definitions. The general classes of choice rules can be too complex to articulate effectively (Echenique (2007)). Choice rules in the GS family appear naturally in important real-world applications. For example, policy goals and affirmative action restrictions in India uniquely identify a particular choice function for reserve categories, the two-step meritorious horizontal choice rules (Sönmez and Yenmez (2022)). We show that an aggregate choice rule of an institution in which each division uses the two-step meritorious horizontal choice rule is in the GS family under any monotone transfer function (Proposition 1). In addition, various choice rules that are being implemented and a new one proposed by Hu and Zhang (2025) for admission to high schools in China are also in the GS family.
Other prominent choice rules in the GS family include the reserve-based choice rules of Echenique and Yenmez (2015), the slot-specific priority choice rules of Kominers and Sönmez (2016), and the dynamic reserves choice rules of Aygün and Turhan (2020) and the multi-price choice rule of Greenberg, Pathak, and Sönmez (2024), among many others. GS choice rules offer considerable latitude to integrate complex diversity constraints. Moreover, the general framework we introduce is flexible to accommodate many other real-world matching applications, such as airline seat upgrades (Kominers and Sönmez (2016)), multidimensional reserves in Brazil (Aygün and Bó (2021)), high school admissions in China (Hu and Zhang (2025)), affirmative action implementations in Chilean school choice (Correa et al. (2021)), among others. Thus, the GS family provides a unified framework for market designers.
The paper is organized as follows. Section 2 discusses the related literature and contributions of our paper. Section 3 presents the model that formulates the general theory of sequential reserves for a single institution. Section 4 formulates the centralized matching problem with multiple institutions. Section 4 introduces the completion of a generalized sequential choice rule and thoroughly investigates its properties. The main result of the paper on stable and strategy-proof mechanisms through GS choice rules is presented in Section 4. Section 5 presents two important real-world applications of the GS choice rules: affirmative action in India and admissions to high schools in China. Finally, Section 6 concludes. All formal proofs are presented in the appendix.
2. Literature Review
This paper offers two important contributions to the literature. First, we develop a generalized theoretical framework for the implementation of soft reserve policies, within a single institution, where reserve categories’ choice rules transcend the selection based on simple baseline priority orderings. Second, we extend our analysis to centralized markets with multiple institutions, demonstrating that both stability and strategy-proofness are preserved when employing the comprehensive family of GS choice rules.
The paper closest to ours is Hatfield, Kominers, and Westkamp (2017) in which the authors study matching markets where institutions have multiple divisions that face capacity constraints similar to ours. The authors also study how hiring in one division affects the resources available for hiring in other divisions. Unlike our GS choice rules, the complexity of choice rules with flexible allotments they introduce could make them challenging to implement in practice. Their allotment functions require determining how many positions are allocated to each division given the available set of alternatives.
This paper contributes to the existing literature on diversity constraints in matching market design. The idea of reserve was introduced in the seminal work by Hafalir, Yenmez, and Yildirim (2013) after Kojima (2012) pointed out the drawbacks of the quota policy. Ehlers, Hafalir, Yenmez, and Yildirim (2014) show that, with hard bounds, there might not exist assignments that satisfy standard fairness and non-wastefulness properties. To achieve fair and non-wasteful assignments, the authors propose to interpret constraints as soft bounds. There is now a rich market design literature on implementing diversity in matching markets. Some other examples are Kamada and Kojima (2015), Biró, Hassidim, Romm, Shorrer, and Sóvágó (2020), and Hassidim, Romm, and Shorrer (2021).
This paper contributes to the growing body of literature on resource allocation problems in India. In Section 5.1, we examine resource allocation problems in India under complex affirmative action constraints as a key application. We show that employing the meritorious horizontal choice rule (Sönmez and Yenmez (2022)) within each division (or reserve category) in conjunction with monotone transfer functions produces choice rules within the GS family. Our formulation explicitly accounts for the de-reservation of vacant Other Backward Classes (OBC) positions, a consideration absent from Sönmez and Yenmez (2022). [^2] Aygün and Turhan (2022) examine the concurrent implementation of vertical reservations and OBC de-reservations, specifically focusing on the recently revamped admissions procedures for technical universities in India, as originally designed and implemented by Baswana, Chakrabarti, Chandran, Kanoria, and Patange (2019). Furthermore, our model allows individuals to express preferences over pairs of institution-position category, improving the applicability of the framework to real-world allocation scenarios. Echenique and Yenmez (2015) is the first paper to present Indian affirmative action as an example of controlled school choice. Subsequent work by Aygün and Turhan (2017) delves into the specific challenges in admissions to the Indian Institutes of Technologies (IITs). Aygün and Turhan (2020) study implementation of vertical reservations and de-reservations in the context of admissions to IITs where the authors introduced a family of dynamic reserves choice rules. The family of GS choice rules presented in the current paper nests these dynamic reserve choice rules. This generalization is of significance, as the dynamic reserve choice rules proposed by Aygün and Turhan (2020) were formulated on the basis of divisions’ q-responsive choice rules, which are inherently limited to incorporate horizontal reservations.
3. A General Theory of Sequential Reserves
Consider an institution $s$ that has a set of divisions $\mathcal{K}_{s}=\{1,...,K_{s}\}$. Each division $k\in\mathcal{K}_{s}$ has an associated choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ that specifies the contracts that division $k$ chooses given a set of alternatives and a capacity that may vary as a function of the number of vacant positions in the preceding divisions. We require that for a set of contracts $Y\subseteq\ X_{s}$ and a capacity of $\kappa\in\mathbb{Z}_{+}$, $\mid C_{s}^{k}(Y;\kappa)\mid\leq\kappa$.
Given a set contracts $Y\equiv Y^{1}\subseteq\ X$ , a capacity $\overline{q}_{s}$ for institution $s$, and a capacity $\overline{q}_{s}^{1}$ for division 1, we compute the set chosen $C_{s}(Y,\overline{q}_{s})$ sequentially in $K_{s}$ steps, where division $k$ chooses in step $k$, for $k=1,...,K_{s}$, as follows:
Step 1
Given $Y^{1}$ and $\overline{q}_{s}^{1}$, division 1 chooses $C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})$. Let
$r_{1}=\overline{q}_{s}^{1}-\mid C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})\mid$
be the number of vacant positions in division 1 and
$Y^{2}\equiv Y^{1}\setminus\{x\in Y^{1}\mid\mathbf{i}(x)\in\mathbf{i}[C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})]\}$
be the set of contracts of individuals who were not selected by division 1.
Step k ($2\leq k\leq K_{s})$
Given the set of remaining contracts $Y^{k}$ and division $k$’s capacity $q_{s}^{k}(r_{1},...,r_{k-1})$, division $k$ chooses $C_{s}^{k}(Y^{k};q_{s}^{k}(r_{1},...,r_{k-1}))$. Let
$r_{k}=q_{s}^{k}(r_{1},...,r_{k-1})-\mid C_{s}^{k}(Y^{k};q_{s}^{k}(r_{1},...,r_{k-1}))\mid$
be the number of unfilled positions in division $k$ and
$Y^{k+1}\equiv Y^{k}\setminus\{x\in Y^{k}\mid\mathbf{i}(x)\in\mathbf{i}[C_{s}^{k}(Y^{k};q_{s}^{k}(r_{1},...,r_{k-1}))]\}$
be the set of contracts of individuals who were not selected by any division in the first $k$ steps.
The union of divisions’ choices is the institution’s chosen set. That is,
$C_{s}(Y,q_{s})\equiv C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})\cup\bigcup_{k=2}^{K_{s}}C_{s}^{k}(Y^{k};q_{s}^{k}(r_{1},...,r_{k-1})).$
Transfer Policy
Given an initial capacity of the first division $\overline{q}_{s}^{1}$, a transfer policy of institution $s$ is a sequence of capacity functions $q_{s}=(\overline{q}_{s}^{1},(q_{s}^{k})_{k=2}^{K_{s}})$, where $q_{s}^{k}:\>\mathbb{Z}_{+}^{k-1}\longrightarrow\mathbb{Z}_{+}$ for all $k\in\mathcal{K}_{s}$ and such that
$\overline{q}_{s}^{1}+q_{s}^{2}(0)+q_{s}^{3}(0,0)+\cdots+q_{s}^{K}(0,...,0)=\overline{q}_{s}.$
A transfer policy $q_{s}$ is $\mathbf{monotonic}$ if, for all $j\in\{2,...,K_{s}\}$ and all pairs of sequences $(r_{l},\widetilde{r}_{l})$, such that $\widetilde{r}_{l}\geq r_{l}$ for all $l\leq j-1$,
$q_{s}^{j}(\widetilde{r}_{1},...,\widetilde{r}_{j-1})\geq q_{s}^{j}(r_{1},...,r_{j-1}),\;\text{and}$
$\sum_{m=2}^{j}[q_{s}^{m}(\tilde{r}_{1},...,\widetilde{r}_{m-1})-q_{s}^{m}(r_{1},...,r_{m-1})]\leq\sum_{m=1}^{j-1}[\widetilde{r}_{m}-r_{m}].$
Conditions on Divisions’ Choice Rules
A choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ satisfies the substitutes property if, for all $q\in\mathbb{Z}_{+}$ and for all $x,y\in X$ and $Y\subseteq X$,
$y\notin C_{s}^{k}(Y\cup\{y\};q)\;\Longrightarrow\;y\notin C_{s}^{k}(Y\cup\{x,y\};q).$
A choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ satisfies size monotonicity if, for all $q\in\mathbb{Z}_{+}$ and for all contracts $x\in X$ and sets of contracts $Y\subseteq X$, we have
$\mid C_{s}^{k}(Y;q)\mid\leq\mid C_{s}^{k}(Y\cup\{x\};q)\mid.$
The substitutes condition and size monotonicity jointly imply the irrelevance of rejected contracts (IRC) condition (Aygün and Sönmez (2013)). A choice rule $C_{s}^{k}$ satisfies the IRC condition if, for any $Y\subseteq X$ and any $x\in X\setminus Y$,
$x\notin C_{s}^{k}(Y\cup\{x\};q)\text{ implies }C_{s}^{k}(Y;q)=C_{s}^{k}(Y\cup\{x\};q).$
Both the substitutes property and size monotonicity are standard conditions that impose restrictions on the choice function when the set of alternatives changes, while the capacity remains unchanged. The next property imposes conditions on the choice function when the capacity increases while the set of alternatives remains unchanged. This property is in the spirit of the expansion property of the choice rules defined in Chambers and Yenmez (2017).
Definition 1.
A choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ satisfies quota monotonicity (QM) if for any $q,\Delta\in\mathbb{\mathbb{Z}}_{+}$ such that for all $Y\subseteq X$,
$C_{s}^{k}(Y,q)\subseteq C_{s}^{k}(Y,q+\Delta),$
and
$\mid C_{s}^{k}(Y,q+\Delta)\mid-\mid C_{s}^{k}(Y,q)\mid\leq\Delta.$
The QM property requires that (1) if there is an increase in capacity, the choice rule selects every contract it was choosing before the increase in capacity from any given set of contracts, and (2) if the capacity increases by $\Delta$, then the difference between the number of contracts chosen with the increase in capacity and the initial capacity cannot exceed $\Delta$. The first condition is the expansion property of Chambers and Yenmez (2017). The second condition, which is similar to the non-excessive reduction property of Westkamp (2013), restricts the magnitude of expansion when capacity is changed.
We are now ready to define a generalized sequential (GS) choice rule.
Definition 2.
An aggregate choice rule $C_{s}$ is an GS rule if it can be represented as a pair $\left(\left(C_{s}^{k}\right)_{k\in\mathcal{K}_{s}},q_{s}\right)$ of a collection of division choice rules and transfer policy, where each division’s choice rule satisfies the substitution condition, the size monotonicity and the quota monotonicity, and the transfer policy is monotone.
The family of GS choice rules is a parametrized family of aggregate choice rules derived from choice functions that satisfy both the substitutes condition and the size monotonicity. This parametrization stems from two primary sources. First, the transfer function provides one dimension of parametrization. Different monotonic transfer functions generate distinct aggregate choice rules within this family. Second, parametrization arises from the expansion of divisions’ choice rules while maintaining quota monotonicity. Within our theoretical framework, a division may be characterized by a sequence of choice rules, with each rule corresponding to a specific integer capacity. [^3] By defining the choice rule of division $k$ as $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ we emphasize that capacity of division $k$ may vary. In a sense, this definition assigns a series of choice functions to division $k$, one for each given integer capacity. When the capacity of a substitutable and size monotonic choice rule is adjusted, the result is a new choice rule that also satisfies these properties. The quota monotonicity property establishes the precise relationship between these capacity-differentiated choice rules. It is important to note that a substitutable and size monotonic choice rule can be "expanded" to another such rule through multiple different ways, further enriching the parametric space of the GS family.
When each division’s choice rule is q-responsive, which implies the substitutes property, size monotonicity, and QM, the aggregate choice rule becomes a dynamic reserves choice rule introduced in Aygün and Turhan (2020). Therefore, the family of GS rules nests the family of dynamic reserves choice rules. An important example of a choice rule that satisfies the substitute property, size monotonicity, and QM but violates responsiveness is the meritorious horizontal choice rule introduced in Sönmez and Yenmez (2022). The GS family also nests the family of slot-specific priorities choice rules, introduced by Kominers and Sönmez (2016), because Aygün and Turhan (2020) shows that each slot-specific priorities choice rule can be represented as a dynamic reserves choice rule.
It is important to note that the aggregate choice rule in the GS family may fail to satisfy the substitute property and size monotonicity, even though the divisions’ choice rules satisfy them. However, a stable and strategy-proof mechanism design is still possible.
4. Stable and Strategy-proof Mechanism Design
We now consider a centralized market in which there is a finite set of individuals $I=\{i_{1},...,i_{n}\}$ and a finite set of institutions $S=\{s_{1},...,s_{m}\}$. We let $\overline{q}_{s}$ be the capacity of the institution $s$. There is a finite set of contracts $X$. $\mathbf{i}(x)$ and $\mathbf{s}(x)$ denote the individual and institution associated with the contract $x\in X$, respectively.
For any set of contracts $Y\subseteq X$, let $\mathbf{i}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{i}(x)\}$ be the set of individuals who have a contract associated with them in $Y$. Similarly, we let $\mathbf{s}(Y)\equiv\bigcup_{x\in Y}\{\mathbf{s}(x)\}$. We also define $Y_{i}\equiv\{x\in Y\mid i=\mathbf{i}(x)\}$ for each $i\in I$ and $Y_{s}\equiv\{x\in Y\mid s=\mathbf{s}(x)\}$ for each $s\in S$.
Each individual $i\in I$ has a strict preference $P_{i}$ over contracts in $X_{i}$ and an outside option $\emptyset$. A contract $x\in X_{i}$ is acceptable to $i$ with respect to $P_{i}$ if $xP_{i}\emptyset$. We let $R_{i}$ be the weak preference relation associated with $P_{i}$. $xR_{i}y$ if and only if $xP_{i}y$ or $x=y$.
Each institution $s\in S$ has multiunit demand and is endowed with an aggregate choice rule $C_{s}$ that describes how $s$ would choose from any set of contracts. For all $Y\subseteq\ X$ and $s\in S$, the aggregate choice rule $C_{s}$ is such that $C_{s}(Y)\subseteq Y_{s}$ and $\mid C_{s}(Y)\mid\leq\overline{q}_{s}$.
A matching is a set of contracts $Y\subseteq X$. A matching $Y\subseteq X$ is feasible if $\mid Y_{i}\mid\leq 1$ for all $i\in I$. A feasible outcome $Y\subseteq X$ is stable if
  1. (1)
    $Y_{i}R_{i}\emptyset$, for all $i\in I$,
  2. (2)
    $C_{s}(Y)=Y_{s}$, for all $s\in S$, and
  3. (3)
    there is no $Z\subseteq(X\setminus Y)$, such that $Z_{s}\subseteq C_{s}(Y\cup Z)$ for all $s\in\mathbf{s}(Z)$ and $Z_{i}P_{i}Y_{i}$ for all $i\in\mathbf{i}(Z)$.
Condition (1) is the standard individual rationality condition for individuals. Condition (2) requires that institutions’ selection procedures are respected. This condition ensures that criteria, such as diversity and affirmative action, that are encoded in institutions’ choice rules are complied. [^4] See Alva and Dogan (2023) for a detailed discussion of this property. Condition (3) is the standard no blocking condition and requires that there be no set of contracts $Z$ such that all institutions and individuals associated with contracts in $Z$ prefer to receive the contracts in $Z$.
Stability and strategy-proofness are fundamental objectives in market design. Stability precludes scenarios where agents can profitably rearrange their matches after market-clearing. Strategy-proofness establishes truthful preference revelation as a dominant strategy; this eliminates incentives for strategic misrepresentation, thereby simplifying market participation and ensuring that allocations derive from accurate preference data.
Given a profile of aggregate choice rules of institutions $C=(C_{s})_{s\in\ S}$, a mechanism $\psi(\cdot;C)$ maps preference profiles $P=(P_{i})_{i\in\ I}$ to outcomes. Unless otherwise stated, we assume that institutions’ choice rules are fixed. A mechanism $\psi$ is stable if $\psi(P;C)$ is a stable outcome for every preference profile $P$, given a profile of the institutions’ choice rules $C$.
A mechanism $\psi$ is strategy-proof if, given a profile of institutions’ choice rules $C$, for every preference profile $P$ and for each individual $i\in I$, there is no $\widetilde{P}_{i}$ such that
$\psi(\widetilde{P}_{i},P_{-i};C)P_{i}\psi(P;C).$
Hatfield and Kominers (2019) introduce conditions on the choice rules of the institutions to guarantee a stable and strategy-proof matching mechanism.
Definition 3 (Hatfield and Kominers (2019)).
A completion of a choice function $C_{s}$ is a choice function $\overline{C}_{s}$, such that for all $Y\subseteq X$, either $\overline{C}_{s}(Y)=C_{s}(Y)$ or there exists a distinct $x,x^{{}^{\prime}}\in\overline{C}_{s}(Y)$ such that $\mathbf{i}(x)=\mathbf{i}(x^{{}^{\prime}})$.
Definition 4 (Hatfield and Kominers (2019)).
If a choice function $C_{s}$ has a completion that satisfies the substitute property and the IRC condition, then $C_{s}$ is said to be substitutably completable. If every choice rule in a profile $C=(C_{s})_{s\in S}$ is substitutably completable, then $C$ is substitutably completable.
Let $C_{s}(\cdot;q_{s})$ be an GS choice rule induced by the transfer function $q_{s}$. We define a related choice function $\overline{C}_{s}(\cdot;q_{s})$ as follows:
Given a set of contracts $Y\equiv Y^{1}\subseteq X$, a capacity $\overline{q}_{s}$ for institution $s$, and a capacity $\overline{q}^{1}_{s}$ for division 1, we compute the set of chosen contracts $\overline{C}_{s}(Y;q_{s})$ in steps $K_{s}$, where divisions choose in sequential order.
Step 1
Given $Y^{1}$ and $\overline{q}^{1}_{s}$, division 1 chooses $\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})$. Let $r_{1}=\overline{q}^{1}_{s}-\mid\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\mid$. Let $Y^{2}\equiv Y^{1}\setminus\overline{C}^{1}_{s}\left(Y^{1};\overline{q}^{1}_{s}\right)$.
Step k. ($2\leq k\leq K_{s})$
Given $Y^{k}$ and its capacity $q^{k}_{s}(r_{1},...,r_{k-1})$, division $k$ chooses $\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))$. Let $r_{k}=q^{k}_{s}(r_{1},...,r_{k-1})-\mid\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))\mid$. Let $Y^{k+1}\equiv Y^{k}\setminus\overline{C}^{k}_{s}\left(Y^{k};q^{k}_{s}\left(r_{1},\ldots,r_{k-1}\right)\right)$.
The union of divisions’ choices is the institution’s chosen set. That is,
$\overline{C}_{s}(Y;q_{s})\equiv\overline{C}^{1}_{s}(Y^{1};\overline{q}^{1}_{s})\cup\bigcup_{k=2}^{K_{s}}{\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1}))}.$
The difference between $C_{s}$ and $\overline{C}_{s}$ is as follows: In the computation of $C_{s}$, if a contract of an individual is chosen by some category, then her other contracts are removed for the rest of the procedure. According to $\overline{C}_{s}$, if an individual’s contract is chosen by division $k$, then her other contracts will still be available for the remaining divisions.
Theorem 1.
$\overline{C}_{s}$ is a completion of $C_{s}$. Moreover, $\overline{C}_{s}$ satisfies the IRC condition, the substitute property, and the size monotonicity.
Cumulative offer mechanisms (COMs) have received special attention in the literature. In an COM, individuals propose contracts according to a strict ordering $\triangleright$ of the elements of $X$. In each step, some individual who does not currently have a contract held by any institution proposes his most preferred contract that has not yet been proposed. Then, each institution chooses its most preferred set of contracts according to its choice rule and holds this set until the next step. When multiple individuals are able to propose in the same step, the individual who proposes is determined by the ordering $\triangleright$.
Given the aggregate choice rules of the institutions and the preferences of individuals, the outcome of the COM is calculated by the cumulative offer process (COP) as follows:
Step 1
Some individual $i^{1}\in\ I$ proposes her most preferred contract, $x^{1}\in\ X_{i^{1}}$. Institution $\mathbf{s}\left(x^{1}\right)$ holds $x^{1}$ if $x^{1}\in C_{\mathbf{s}\left(x^{1}\right)}\left(\left\{x^{1}\right\}\right)$, and rejects $x^{1}$ otherwise. Set $A_{\mathbf{s}\left(x^{1}\right)}^{2}=\left\{x^{1}\right\}$, and set $A_{s^{\prime}}^{2}=\emptyset$ for each $s^{\prime}\neq\mathbf{s}\left(x^{1}\right)$, i.e., the sets of contracts available to institutions at Step 2.
Step l
Some individual $i^{l}\in\ I$, for whom no institution currently holds a contract, proposes her most preferred contact that has not yet been rejected $x^{l}\in X_{i^{l}}\setminus\left(\bigcup_{j\in S}A_{j}^{l}\right)$. Institution $\mathbf{s}\left(x^{l}\right)$ holds the set of contracts in $C_{\mathbf{s}\left(x^{l}\right)}\left(A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}\right)$ and rejects all other contracts in $A_{s\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$; institutions $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$ continue to hold all contracts they held at the end of Step $l-$1. Set $A_{\mathbf{s}\left(x^{l}\right)}^{l+1}=A_{\mathbf{s}\left(x^{l}\right)}^{l}\cup\left\{x^{l}\right\}$, and set $A_{s^{\prime}}^{l+1}=A_{s^{\prime}}^{l}$ for each $s^{\prime}\neq\mathbf{s}\left(x^{l}\right)$.
The algorithm terminates when no individual can propose. Each institution is assigned the set of contracts that it holds in the last step.
Given $P=\left(P_{i}\right)_{i\in I}$ and $C=\left(C_{s}\right)_{s\in S}$, let $\Phi\left(P,C\right)$ be the result of the COM. Let $\Phi_{i}\left(P,C\right)$ and $\Phi_{s}\left(P,C\right)$ denote the assignment of individual $i$ and institution $s$, respectively.
A direct corollary of Theorem 1 is that the COM is the only candidate if the policy makers insist on stability and strategy-proofness.
Corollary 1.
The COM is the unique mechanism that is stable and strategy-proof when institutions’ choice rules are in the GS family.
5. Practical Choice Rules in the GS Family
Policy goals in many real-world matching markets naturally give rise to choice rules that are in the GS family. In this section, we present three applications of real-world resource allocation problems where the proposed choice rules that emerge from policy desiderata are in the GS family. Namely, we analyze (i) the aggregate choice rule in which the meritorious horizontal choice rule selects candidates in each category in allocation government jobs in India (Sönmez and Yenmez (2022) and (ii) the simultaneous open-reserve open-choice rules proposed for high school admissions in China (Hu and Zhang (2025)).
5.1. Affirmative Action in India
India has implemented a complex affirmative action program in allocating government jobs and admissions to public universities since the 1950s. India’s highly regulated reserve system consists of two different provisions: vertical and horizontal. [^5] The terms vertical and horizontal first appeared in the Supreme Court judgment in Indra Sawhney vs. Union of India (1992). The case is available at https://indiankanoon.org/doc/1363234/.
Vertical Reservations
The reservations for Scheduled Castes (SC), Scheduled Tribes (ST), Other Backward Classes (OBC) and Economically Weaker Sections (EWS) are called vertical reservations. 15%, 7.5%, 27%, and 10% of the positions available at institutions are designated for these respective groups. [^6] In 2019, the Indian legislative body instituted a 10% vertical reservation for a subset of the general category, specifically the Economically Weaker Section (EWS), defined by an annual income threshold below Rs. 8 lakhs. Individuals not included in these vertical categories are classified as members of the general category (g). The residual 40.5% of available positions is referred to as open category positions. Any open category positions secured by individuals belonging to SC, ST, OBC, or EWS are not to be deducted from their respective vertical reservations by law. Failure to declare membership in the SC, ST, OBC, or EWS categories relegates candidates to category-g by default. Importantly, disclosure of vertical category affiliation is discretionary.
Let $R=\{SC,ST,OBC,EWS\}$ be the set of reserved categories. In each institution $s\in S$, $\overline{q}_{s}^{SC}$, $\overline{q}_{s}^{ST}$, $\overline{q}_{s}^{OBC}$, and $\overline{q}_{s}^{EWS}$ positions are earmarked for the categories SC, ST, OBC, and EWS, respectively. The remaining $\overline{q}_{s}^{o}=\overline{q}_{s}-\left(\overline{q}_{s}^{SC}+\overline{q}_{s}^{ST}+\overline{q}_{s}^{OBC}+\overline{q}_{s}^{EWS}\right)$ positions are the open category positions. We denote the set of vertical categories $V=\{o,SC,ST,OBC,EWS\}$, where $o$ denotes the open category. Let
$\overline{q}_{s}=(\overline{q}_{s}^{o},\overline{q}_{s}^{SC},\overline{q}_{s}^{ST},\overline{q}_{s}^{OBC},\overline{q}_{s}^{EWS})$
be the vector of vertical reservations at institution $s$.
The function $t:I\rightarrow R\cup\left\{g\right\}$ denotes individuals’ category membership. For every individual $i\in I$, $t(i)$ denotes the category to which individual $i$ belongs. We denote a reserved category membership profile by $T=\left(t(i)\right)_{i\in I}$, and let $\mathcal{T}$ be the set of all possible reserved category membership profiles.
Horizontal Reservations
Horizontal reservations are provided to other marginalized groups, such as women and individuals with disabilities [^7] For an in-depth analysis, see Sönmez and Yenmez (2022).. The SCI’s judgment in Indra Sawhney (1992) mandates that horizontal reservations must be implemented within each vertical category on a minimum guarantee basis. That is, in each vertical category, horizontally reserved positions must be allocated before any unreserved positions. However, the SCI judgments have left it ambiguous whether an individual eligible for multiple types of horizontal reservation is counted against a single horizontal category or multiple categories concurrently.
Let $H=\{h_{1},\ldots,h_{L}\}$ be the set of horizontal types. The correspondence $\rho:\>I\rightrightarrows H\cup\{\emptyset\}$ represents the horizontal types of individuals. $\rho(i)\subseteq H\cup\{\emptyset\}$ denotes the set of horizontal types that the individual $i$ can claim. $\rho(i)=\emptyset$ means that $i$ has no horizontal type. We denote by $\kappa_{v}^{j}$ the number of positions reserved for horizontal type $h_{j}\in H$ in the vertical category $v\in V$. The vector $\kappa_{v}^{s}\equiv(\kappa_{v}^{1},...,\kappa_{v}^{L})$ denotes the horizontal reservations in the vertical category $v\in V$ in the institution $s$. Let $\kappa_{s}\equiv\left\{\kappa_{v}^{s}\right\}_{v\in V}$ denote the horizontal reservations of the institution $s$.
De-reservations
SC and ST are classified as hard reserves, which means that the vacant positions reserved for these groups cannot be transferred to other categories. Although reservations for OBC are treated as hard reserves in the context of governmental job allocations, they were suggested and have been implemented as soft reserves in admissions to public colleges, pursuant to the landmark SCI decision in Ashoka Kumar Thakur vs. Union of India (2008). [^8] The judgment is available at https://indiankanoon.org/doc/1219385/. The SCI rulings did not provide an explicit procedural framework for de-reserving unfilled OBC positions. The lack of clear rules has led to various inconsistent practices in school admissions in India (Aygün and Turhan (2022)).
Priority Orders
Each institution $s$ has a strict priority order $\succ_{s}$ over ${}\mathcal{I}\cup\{\emptyset\}$. We write $i\succ_{s}j$ to mean that the applicant $i$ has a higher priority than $j$ at $s$. Similarly, we write $i\succ_{s}\emptyset$ to say that applicant $i$ is acceptable for $s$. $\emptyset\succ_{s}i$ means that the applicant $i$ is unacceptable for $s$. The profile of institutions’ priorities is denoted $\succ=(\succ_{s_{1}},...,\succ_{s_{m}})$. For each institution $s\in\mathcal{S}$, the merit ordering for individuals of type $r\in\mathcal{R}$, denoted by $\succ_{s}^{r}$, is obtained from $\succ_{s}$ in a straightforward manner as follows:
  • for $i,j\in\mathcal{I}$ such that $t_{i}=r$, $t_{j}\neq r$, $i\succ_{s}\emptyset$, and $j\succ_{s}\emptyset$, we have $i\succ_{s}^{r}\emptyset\succ_{s}^{r}j$, where $\emptyset\succ_{t}^{r}j$ means individual $j$ is unacceptable for category $r$ at institution $s$.
  • for any other $i,j\in\mathcal{I}$, $i\succ_{s}^{r}j$ if and only if $i\succ_{s}j$.
Individuals’ Preferences
We define $\mathcal{X}\equiv\underset{t(i)=g}{\bigcup}\{i\}\times S\times\{o\}\underset{t(i)\in R}{\bigcup}\{i\}\times S\times\{r,o\}$ as the set of all contracts. Each contract $x\in\mathcal{X}$ is between an individual $\mathbf{i}(x)$ and an institution $\mathbf{s}(x)$ and specifies a vertical category $\mathbf{t}(x)$ in which the individual $\mathbf{i}(x)$ is admitted.
Each individual in category g has a preference for $\mathcal{S\times}\left\{o\right\}\cup\left\{\emptyset\right\}$. Each member of the reserve category $r$ has a preference over $\mathcal{S}\times\left\{r,o\right\}\cup\left\{\emptyset\right\}$, where $\emptyset$ denotes the outside option. [^9] We assume that individuals do not have preferences for horizontal types as there is no evidence that individuals care about the horizontal types under which they are admitted. All discussion on reservation policy in India centers on vertical reservations. We write $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ to mean that the individual $i$ strictly prefers admission to institution $s$ through the vertical category $v$ to admission to institution $s^{{}^{\prime}}$ via vertical category $v^{{}^{\prime}}$. The at-least-as-well relation $R_{i}$ is obtained from $P_{i}$ as follows: $(s,v)R_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ if and only if $(s,v)P_{i}(s^{{}^{\prime}},v^{{}^{\prime}})$ or $(s,v)=(s^{{}^{\prime}},v^{{}^{\prime}})$. An institution and vertical category pair $(s,v)$ is acceptable to individual $i$ if it is at least as good as the outside option $\emptyset$ and is unacceptable to her if it is worse than the outside option $\emptyset$.
Meritorious Horizontal Choice Rule
We consider one-to-one horizontal matching, where an admitted applicant with multiple eligible horizontal types is counted against only one of these types. Sönmez and Yenmez (2022) proposed the meritorious horizontal choice rule, denoted $C^{m}$, in this setting. Within each vertical category $v$, the authors consider the assignment of horizontally reserved positions $H^{v}=\bigcup_{j=1}^{L}H_{j}^{v}$ to a set of individuals $I$ as a one-to-one matching problem, where $H_{j}^{v}$ denotes the set of positions reserved for horizontal type $h_{j}\in H$ in vertical category $v$. The number of positions in $H_{j}^{v}$ is denoted by $\kappa_{v}^{j}$.
Given a vertical category $v\in\mathcal{V}$ and a set of individuals $I$ such that $t_{i}=v$ for all $i\in I$, define the one-to-one matching between the individuals and horizontally reserved positions as a mapping $\mu:I\rightarrow H^{v}\bigcup\left\{\emptyset\right\}$ such that for any $i\in I$ and $h_{j}\in H$,
  • $\mu(i)\in H_{j}^{v}$ implies $h_{j}\in\rho(i)$, and for any $i,k\in I$,
  • $\mu(i)=\mu(k)\neq\emptyset$ implies $i=k$.
Let $n^{v}(I)$ represent the maximum number of horizontally reserved positions within the vertical category $v$. The matching between individuals and horizontally reserved positions is said to have “maximum cardinality" if there is no alternative assignment that would strictly increase the total number of positions allocated to eligible individuals. For a given vertical category $v$ and a set of individuals $I$ belonging to category $v$, an individual $i$ is said to “increase horizontal reservation utilization" if $n^{v}(I\cup\{i\})=n^{v}(I)+1$.
Given a vertical category $v$, a set of applicants $A$ such that $\mathbf{v}(i)=v$ for all $i\in A$, a profile of the horizontal types of applicants $\left(\rho(i)\right)_{i\in A}$, the meritorious horizontal choice rule $C^{m}$ selects the applicants as follows:
[…]
Choose the highest merit score individual in $A$ with a horizontal type. Denote this individual by $i_{1}$ and let $A_{1}=\left\{i_{1}\right\}$. If no such individual exists, proceed to Step 2.
[…]
Choose the highest merit score individual in $A\setminus A_{n-1}$ who increase the horizontal reservation utilization given $A_{n-1}$, if any. Denote this individual by $i_{n}$ and let $A_{n}=A_{n-1}\cup\{i_{n}\}$. If no such individual exists, proceed to Step 2.
[…]
Fill remaining positions following the merit score with individuals among the unassigned individuals until all positions are assigned or all individuals are selected.
Fleiner (2001) showed that the first stage of $C^{m}$ satisfies the substitute property. Sönmez and Yenmez (2022) proved that $C^{m}$ satisfies size monotonicity. We prove that $C^{m}$ also satisfies the quota monotonicity. Thus, any aggregate choice rule in which the divisions select according to $C^{m}$ under a monotone transfer policy is in the GS family.
Aggregate Choice Rule with De-reservation
According to the aggregate choice rule with transfer, $C^{mT}$, vacant OBC positions are reverted to an open category and filled following merit scores. The rule $C^{mT}$ is an GS choice rule and can be written as
$C_{s}^{mT}\equiv\left(\left(C_{v}^{m}\right)_{v\in\mathcal{V}},q_{s}\right),$
where
  • the precedence sequence is $o-SC-ST-OBC-EWS-D$, where $D$ denotes “de-reserved positions” that are reverted from OBC, and
  • $q_{s}=\left(q_{s}^{o},q_{s}^{SC},q_{s}^{ST},q_{s}^{OBC},q_{s}^{EWS},q_{s}^{D}\right)$ such that $q_{s}^{v}=\overline{q}_{s}^{v}$ for all $v\in\left\{o,SC,ST,OBC,EWS\right\}$, and $q_{s}^{D}=r_{OBC}$ is the number of vacant OBC positions.
When each vertical category $v\in\mathcal{V}$ chooses applicants through $C_{v}^{m}$, under any monotonic transfer policy, the aggregate choice rules of the institutions are in the GS family.
Proposition 1.
$C_{s}^{mT}$ belongs to the GS family.
[…]
5.2. Affirmative Action in Chinese High School Admissions
In China, public high schools are subject to an affirmative action policy in which high schools reserve a proportion of their available seats for graduates of low-performing middle schools. High school admissions are merit-based and highly competitive. The affirmative action policy requires high schools to reserve 50% or more of their seats exogenously to middle schools.
Hu and Zhang (2025) study a novel aspect of the Chinese affirmative action policy, called the “verifiability," that has been implemented in almost every city in China. Verifiability of an allocation mechanism ensures that students can verify their assignments using only their private information (e.g., their preferences over schools) and publicly observed cutoff scores. It is an important aspect of the transparency and accountability of the admission process. The authors analyze several choice rules that are used in Chinese cities. To describe them, we first introduce their notation.
There is a finite set of students $\mathcal{I}$. Let $q_{s}$ be the capacity of the high school $s$. Let $\mathcal{M}$ be the finite set of middle schools. The high school $s$ reserves $q_{s}^{m}$ seats for students from the middle school $m$. We assume that $\sum_{m\in M}q_{s}^{m}<q_{s}$, where $q_{s}^{o}=q_{s}-\sum_{m\in M}q_{s}^{m}$ denotes the number of open seats in high school $s$. The function $\tau:\mathcal{I}\rightarrow\mathcal{M}$ is such that $\tau(i)$ denotes the middle school student $i$ belongs. Let $\mathcal{I}^{m}=\tau^{-1}(m)$ and $I^{m}\equiv I\cap\mathcal{I}^{m}$.
There are two deferred acceptance-based mechanisms that allocate open and reserved seats simultaneously. These mechanisms use the choice rules $\mathcal{C}^{SimRO}$ and $\mathcal{C}^{SimORO}$ we define below.
5.2.1. The Simultaneous-reserve-open Choice Rule $\mathcal{C}^{SimRO}$
This choice rule is used in cities such as Guiyan, Qingdao, and Tianjin for high school admissions in China.
For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimRO}$ selects the students as follows:
Step 1
The reserve seats are allocated for each $m\in\mathcal{M}$ to the highest priority students from $I^{m}$ to capacity $q_{s}^{m}$.
Step 2
All remaining seats are assigned as open seats to the highest ranked students among the remaining until all seats are allocated or all students are admitted.
Note that unfilled reserved seats are allocated as open seats in the second stage. The choice rule $\mathcal{C}^{SimRO}$ is in the GS family.
5.2.2. The Simultaneous-open-reserve-open Choice Rule $\mathcal{C}^{SimORO}$
In Chinese cities Taiyuan, Zengzhou, and Hefei, among many others, for high school admissions, the deferred acceptance mechanism with the choice rule $\mathcal{C}^{SimORO}$ is used.
For any given set $I\subseteq\mathcal{I}$ of students, the choice rule $\mathcal{C}^{SimORO}$ selects the students in three steps.
Step 1
Open seats are allocated to students with the highest priority, regardless of their type, up to capacity $q_{s}^{o}$.
Step 2
The reserve seats are allocated for each $m\in\mathcal{M}$ to the remaining students with the highest priority from $I^{m}$ up to capacity $q_{s}^{m}$.
Step 3
All remaining seats are assigned as open seats to the highest ranked students among the remaining until all seats are allocated or all students are admitted.
The unfilled reserved seats in the second stage are allocated as open seats in the third stage. The choice rule $\mathcal{C}^{SimORO}$ is in the GS family.
Deferred Acceptance Induced Choice Rules
Another version of the deferred acceptance mechanism used for high school admissions in some Chinese cities differs from the two mechanisms above in that each school is divided into two subschools, open subschool and reserved subschools. For school $s\in\mathcal{S}$, the open subschool $s^{o}$ has capacity $q_{S}^{o}$, and the reserved subschool $S^{r}_{s}$ has the capacity vector $q_{s}^{r}=(q_{S}^{m})_{m\in\mathcal{M}}$. We instead model it in matching with contracts framework.
Students are required to submit a preference list of subschools. Then, the student-proposing DA coupled with the following choice rules is run to find a matching. In each round of the DA, each unassigned student applies to the highest-ranked subschool in their preference lists.
We denote a contract between the student $i$ and the school $s$ by $x=(i,s,t)$, where $t\in\{o,r\}$ denotes the seat category under which the student is accepted. Since each student belongs to a middle school in $\mathcal{M}$, each student has two contracts with a given high school. Let $\mathcal{X}=\mathcal{I}\times\mathcal{S}\times\{o,r\}$ denote the set of all contracts. For a given set of contracts $X\in\mathcal{X}$, we let $x_{i}$ and $X_{s}$ denote the set of contracts associated with the student $i$ and the school $s$, respectively. Similarly, let $X_{o}$ and $X_{r}$ denote the contracts with terms $o$ and $r$, respectively. Note that $X_{o}\cup X_{r}=X$.
The Simultaneous-separate Choice Rule $\mathcal{C}^{SimSep}$
Given a set of contracts $X$, the choice rule $\mathcal{C}^{SimSep}$ selects contracts as follows:
Step 1
Consider only the contracts in $X_{o}$. Choose contracts associated with the highest-ranked students up to capacity $q_{s}^{o}$.
Step 2
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students in $(X_{r})_{i}\cap I^{m}$ up to capacity $q_{s}^{m}$.
This rule does not allocate reserved positions that are left unfilled. Thus, it is wasteful. Note that $\mathcal{C}^{SimSep}$ is in the GS family.
The Simultaneous-flexible Choice Rule $\mathcal{C}^{SimFlex}$
Given a set of contracts $X$, the choice rule $\mathcal{C}^{SimFlex}$ selects contracts as follows:
Step 1
Consider only the contracts in $X_{r}$. For each $m\in\mathcal{M}$, choose contracts of the highest-ranked students in $(X_{r})_{i}\cap I^{m}$ up to capacity $q_{s}^{m}$.
Step 2
Consider only the contracts in $X_{o}$. Allocate all remaining seats (open seats and leftover reserved seats) by choosing contracts associated with the highest-ranked students.
$\mathcal{C}^{SimFlex}$ is a GS choice rule.
The Simultaneous-open-reserve Choice Rule $\mathcal{C}^{SimOR}$
Let $I$ be the set of students and $s$ be the only school. Define $X=I\times\{s\}\times\{o,r\}$. Assume that each student $i\in I$ ranks the contract with the open seat over the contract with the reserved seat. Run the student proposing DA while the school $s$ selects students using $\mathcal{C}^{SimFlex}$. The result of the DA procedure defines $\mathcal{C}^{SimFlex}(I)$.
Hu and Zhang (2025) show that $\mathcal{C}^{SimOR}$ does not belong to the class of slot-specific priorities choice rules of Kominers and Sönmez (2016). On the other hand, the GS family contains $\mathcal{C}^{SimOR}$ naturally because $\mathcal{C}^{SimOR}$ satisfies the substitutes condition and is $q$-acceptant, which implies size monotonicity. [^10] Since $\mathcal{C}^{SimOR}$ satisfies the substitutes condition and size monotonicity, it can be considered as a single division’s choice rule.
6. Conclusion
This paper introduced a new family of choice rules called Generalized Sequential (GS) choice rules for institutions with multiple divisions. The GS framework unifies and extends various choice rules that appear naturally in real-world market design applications, particularly those involving complex affirmative action policies. We showed that when institutions use GS choice rules, the Cumulative Offer Mechanism (COM) is the unique mechanism that satisfies both stability and strategy-proofness. This uniqueness result provides strong theoretical support for using COM in settings with complex distributional constraints.
We demonstrated that prominent choice rules used in practice, such as aggregate choice rules that use meritorious horizontal choice rules in each vertical category in India and several choice rules used in high school admissions in numerous cities in China, are members of the GS family. We provide a unified theoretical framework for analyzing them.
The framework developed in this paper provides a foundation for analyzing complex matching markets with distributional constraints. By unifying various existing approaches under the GS umbrella, it opens new avenues for both theoretical analysis and practical market design.
7. Appendix
Proof of Theorem 1
Consider an offer set $Y=Y^{1}\subseteq X$ and assume that there is no pair of contracts $z,z^{{}^{\prime}}\in Y^{1}$ such that $i(z)=i(z^{{}^{\prime}})$ and $z,z^{{}^{\prime}}\in\overline{C}_{s}(Y;q_{s})$. We want to show that $\overline{C}_{s}(Y;q_{s})=C_{s}(Y;q_{s}).$
Let $(Y_{j},r_{j},Y^{j+1})$ and $(Z_{j},\overline{r}_{j},Z^{j+1})$ be sequences of a set of contracts chosen by division $j$, the number of vacant slots in division $j$, and the set of contracts that remains in the choice procedure after division $j$ selects under $C_{s}$ and $\overline{C}_{s}$, respectively.
Given $\bar{q}_{s}^{1}$ and $Y^{1}=Z^{1}$, we have $Z_{1}=\overline{C}^{1}_{s}(Z^{1};\overline{q}_{1}^{s})=C^{1}_{s}(Y^{1};\bar{q}_{s}^{1})=Y_{1}$. Moreover, $\overline{r}_{1}=r_{1}$ and $\overline{q}_{2}^{s}(\overline{r}_{1})=q_{2}^{s}(r_{1})$. Suppose that for all $j\in\{2,...,k-1\}$, we have $Y_{j}=Z_{j}$. We need to show that $Y_{k}=Z_{k}$.
Since $Y_{j}=Z_{j}$ for all $j=1,\ldots,k-1$, we obtain $r_{j}=\overline{r}_{j}$ for all $j=1,\ldots,k-1$. Then $q^{k}_{s}(r_{1},...,r_{k-1})=\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}).$ Since there are no two contracts of an individual chosen by $\overline{C}_{s}$, all the remaining contracts of individuals—whose contracts were chosen by previous divisions—are rejected by $C^{k}_{s}(\overline{Y}^{k};\overline{q}_{s}^{k}(\overline{r}_{1},...,\overline{r}_{k-1}))$. Therefore, the IRC of $C^{k}_{s}$ implies that $C^{k}_{s}(\overline{Y}^{k};\overline{q}_{k}^{s}(\overline{r}_{1},...,\overline{r}_{k-1}))=C_{s}^{k}(Y^{k};q^{k}_{s}(r_{1},...,r_{k-1})).$ Thus, we obtain $\overline{Y}_{k}=Y_{k}$, $\overline{r}_{k}=r_{k}$, and $\overline{q}^{k+1}_{s}(\overline{r}_{1},...,\overline{r}_{k})=q^{k+1}_{s}(r_{1},...,r_{k})$. Since each division selects the same set of contracts under $C_{s}$ and $\overline{C}_{s}$, the result follows.
IRC Condition
Consider $Y=Y^{1}\subseteq X$ such that $Y\neq\overline{C}_{s}(Y;q_{s})$. Take $x\in Y\setminus\overline{C}_{s}(Y;q_{s})$. We need to show that $\overline{C}_{s}(Y;q_{s})=\overline{C}_{s}(Y\setminus\{x\};q_{s}).$
Let $Z^{1}=Y^{1}\setminus\{x\}$. Let $(Y_{j},r_{j},Y^{j+1})$ and $(Z_{j},\tilde{r}_{j},Z^{j+1})$ be sequences of the set of chosen contracts, the number of vacant slots, and the remaining set of contracts for division $j=1,...,K_{s}$ under $\overline{C}_{s}$ from $Y$ and $Z$, respectively.
Since the choice function of division 1 satisfies the IRC, we have $Y_{1}=Z_{1}$. Moreover, $r_{1}=\tilde{r}_{1}$ and $Y^{2}\setminus\{x\}=Z^{2}$. By induction, for each $j=2,...,k-1$, assume that
$Y_{j}=Z_{j},\;r_{j}=\tilde{r}_{j},\;and\;Y^{j}\setminus\{x\}=Z^{j}.$
We need to show that the above equalities hold for $j=k$. Since $x\notin\overline{C}_{s}(Y;q_{s})$ and divisions’ choice functions satisfy the IRC condition, we obtain
$\overline{C}^{k}_{s}(Y^{k};q^{k}_{s}(\overline{r}_{1},...,\overline{r}_{k-1}))=\overline{C}^{k}_{s}(Z^{k};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})).$
By our inductive assumption that $r_{j}=\tilde{r}_{j}$ for each $j=2,...,k-1$, division $k$’s capacity is the same under both choice processes. The number of remaining slots is the same as well. That is, $r_{k}=\tilde{r}_{k}$.
Finally, we know that $x$ is not chosen from the set $Z^{k}\cup\{x\}$, then we have $Y^{k+1}=Z^{k+1}\cup\{x\}.$
Since for all $j\in\{1,...,K_{s}\}$, $Y_{j}=Z_{j}$, we have $\overline{C}_{s}(Y;q_{s})=\overline{C}_{s}(Z;q_{s})$. Thus, $\overline{C}_{s}$ satisfies the IRC.
Substitutes Property
Consider $Y\subseteq X$ such that $Y\neq\overline{C}_{s}(Y;q_{s})$. Let $x$ be such that $x\in Y\setminus\overline{C}_{s}(Y;q_{s})$, and let $z$ be an arbitrary contract in $X\setminus Y$. Let $Z=Y\cup\{z\}$. We need to show that $x\notin\overline{C}_{s}(Z;q_{s}).$
Let $(Y_{j},r_{j},Y^{j})$ and $(Z_{j},\tilde{r}_{j},Z^{j})$ be sequences of the set chosen contracts, the number of vacant slots, and the set of remaining contracts of the divisions $j=1,...,K_{s}$ from $Y$ and $Z$, respectively, under $\overline{C}$.
If $z\in Z\setminus\overline{C}_{s}(Z;q_{s})$, then the IRC of $\overline{C}_{s}$ implies $\overline{C}_{s}(Z;q_{s})=\overline{C}_{s}(Y;q_{s})$. Therefore, $x\notin\overline{C}_{s}(Z;q_{s})$.
Suppose $z\in\overline{C}_{s}(Z;q_{s})$. Consider the first division. If $z\notin Z_{1}$, then by IRC we have $Y_{1}=Z_{1}$ and, thus, $r_{1}=\widetilde{r}_{1}$. Moreover, we obtain $Z^{2}=Y^{2}\cup\{z\}$ and $q^{2}_{s}(r_{1})=q^{2}_{s}(\widetilde{r}_{1})$. Now, consider the case where $z\in Z_{1}$. Note that, by the substitutes condition, all contracts that are rejected from $Y^{1}$ are rejected from $Z^{1}$. By the size monotonicity condition, $\mid Z_{1}\mid\geq\mid Y_{1}\mid$. There are two possibilities:
  1. (1)
    $Z_{1}=Y_{1}\cup\{z\}$, or
  2. (2)
    $Z_{1}=Y_{1}\cup\{z\}\setminus\{w\}$ for some $w\in Y_{1}$.
In the former case, we obtain $Y^{2}=Z^{2}$ and $r_{1}=1+\widetilde{r}_{1}$. Moreover, we have $q^{2}_{s}(\widetilde{r}_{1})\leq q^{2}_{s}(r_{1})\leq 1+q^{2}_{s}(\widetilde{r}_{1})$ due to the monotonicity of capacity transfers. In the latter case, we obtain $Z^{2}=Y^{2}\cup\{w\}$ and $r_{1}=\widetilde{r}_{1}$. Therefore, we have $q^{2}_{s}(r_{1})=q^{2}_{s}(\widetilde{r}_{1})$.
Suppose now that for all $\gamma=j,...,k-1$ we have that either
$\left[Z^{\gamma}=Y^{\gamma}\cup\{w\}\;for\;some\;w\;and\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})=q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\right]$ (1)
or
$\left[Z^{\gamma}=Y^{\gamma}\;and\;q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\leq q^{\gamma+1}_{s}(r_{1},...,r_{\gamma})\leq 1+q^{\gamma+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{\gamma})\right].$ (2)
We have already shown that it holds for $\gamma=1$ and next we show that it also holds for $\gamma=k$.
[…]
$q^{k+1}_{s}(r_{1},...,r_{j})=q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{j}).$
If $w$ is chosen from $Z^{k-1}$, we have two possibilities, depending on whether the capacity of division $k$ is exhausted in the choice process starting with $Y^{0}$ or not. If it is not exhausted, then we obtain
$C_{s}^{k}(Z^{k-1};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))=\{w\}\cup C_{s}^{k}(Y^{k-1};q^{k}_{s}(r_{1},...,r_{k-1})),$
which implies that $Z^{k}=Y^{k}$. Moreover, we have $r_{k}=\tilde{r}_{k}+1$.
The monotonicity of the transfer function implies
$q_{s}^{k+1}(\tilde{r}_{1},...,\tilde{r}_{k})\leq q_{s}^{k+1}(r_{1},...,r_{k})\leq 1+q_{s}^{k+1}(\tilde{r}_{1},...,\tilde{r}_{k}).$
Note that the first inequality follows from the fact that $\tilde{r}_{i}\leq r_{i}$ for all $i=1,...,k$. The second inequality follows from the second condition of the monotonicity of the transfer function.
[…]
We now analyze the latter case, in which we have $Z^{k-1}=Y^{k-1}$ and either $q^{k}_{s}(r_{1},...,r_{k-1})=q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$ or $q^{k}_{s}(r_{1},...,r_{k-1})=1+q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. If $q_{s}^{k}(r_{1},...,r_{k-1})=q_{s}^{k}(\tilde{r}_{1},...,\tilde{r}_{k-1})$, then, given that $Z^{k-1}=Y^{k-1}$, we have $Z^{k}=Y^{k}$. This also implies $r_{k}=\tilde{r}_{k}$. Moreover, we obtain $q^{k+1}_{s}(r_{1},...,r_{k})=q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$ by the monotonicity of the transfer function. Note that $\tilde{r}_{i}\leq r_{i}$ for all $i=1,...,k$ implies $q^{k+1}_{s}(r_{1},...,r_{k})\geq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$ by the first monotonicity condition. The second condition of monotonicity implies $q^{k+1}_{s}(r_{1},...,r_{k})\leq q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})$.
If $q^{k}_{s}(r_{1},...,r_{k-1})=1+q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$, then, given $Z^{k-1}=Y^{k-1}$, we have two cases to consider.
Case 1: If $C_{s}^{k}(Z^{k-1};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))=C_{s}^{k}(Y^{k-1};q^{k}_{s}(r_{1},...,r_{k-1})),$ then we have $Z^{k}=Y^{k}$. In addition, the monotonicity of transfer function implies
$q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k})\leq q^{k+1}_{s}(r_{1},...,r_{k})\leq 1+q^{k+1}_{s}(\tilde{r}_{1},...,\tilde{r}_{k}).$
Case 2: If $C_{s}^{k}(Z^{k-1};q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))\cup\{\vartheta\}=C_{s}^{k}(Y^{k-1},q^{k}_{s}(r_{1},...,r_{k-1}))$ for some $\vartheta$, then we have $Z^{k}=Y^{k}\cup\{\vartheta\}$. Moreover, the monotonicity of the transfer function implies
$q_{s}^{k+1}(r_{1},...,r_{k})=q_{s}^{k+1}(\tilde{r}_{1},...,\tilde{r}_{k}).$
Since $x\notin Y_{k}$, we have $x\notin Z_{k}$ for all $k=1,...,K_{s}$. We conclude that $x\notin\overline{C}_{s}(Y\cup\{z\};q_{s})$.
Size Monotonicity
Take $Y$ and $Z$ such that $Y\subseteq Z\subseteq X$. Let $q_{s}$ be the transfer function. We want to show that
$\mid\overline{C}_{s}(Y;q_{s})\mid\leq\mid\overline{C}_{s}(Z;q_{s})\mid.$
Let $(Y_{j},r_{j},Y^{j})$ and $(Z_{j},\tilde{r}_{j},Z^{j})$ be the sequences of the set of chosen contracts, the numbers of unfilled slots, and the sets of remaining contracts for divisions $j=1,...,K_{s}$ in choice processes starting with $Y=Y^{0}$ and $Z=Z^{0}$, respectively.
For the first division with capacity $\overline{q}_{s}^{1}$, since the choice function is size monotonic, we have
$\mid Y_{1}\mid=\mid C^{1}_{s}(Y^{0};\overline{q}_{s}^{1})\mid\leq\mid C^{1}_{s}(Z^{0};\overline{q}_{s}^{1})\mid=\mid Z_{1}\mid.$
This, in turn, implies $r_{1}=\overline{q}_{s}^{1}-\mid Y_{1}\mid\geq\tilde{r}_{1}=\overline{q}_{s}^{1}-\mid Z_{1}\mid$. Moreover, we have $Y^{1}\subseteq Z^{1}$. To see this, consider $y\in Y^{1}$. This means that $y\notin Y_{1}$. If $y$ is not chosen from $Y^{0}$, then it cannot be chosen from a larger set $Z^{0}$ due to the substitutes property.
Suppose that $\tilde{r}_{j}\leq r_{j}$ and $Y^{j}\subseteq Z^{j}$ hold for all $j=1,...,k-1$. We need to show that
  1. (1)
    $\tilde{r}_{k}\leq r_{k}$, and
  2. (2)
    $Y^{k}\subseteq Z^{k}$.
We first show (1). Given that $\tilde{r}_{j}\leq r_{j}$ for all $j=1,...,k-1$, the first condition of monotonicity implies that $q^{k}_{s}(r_{1},...,r_{k-1})\geq q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. By the size monotonicity, we have
$\mid Z_{k}\mid=\mid C_{s}^{k}(Z^{k-1},q_{s}^{k}(\tilde{r}_{1},...,\tilde{r}_{k-1}))\mid\geq\mid C_{s}^{k}(Y^{k-1},q_{s}^{k}(\tilde{r}_{1},...,\tilde{r}_{k-1}))\mid.$
Moreover, by quota monotonicity, we obtain
$\mid Y_{k}\mid-\mid C_{s}^{k}(Y^{k-1},q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))\mid\leq q^{k}_{s}(r_{1},...,r_{k-1})-q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}).$
The difference on the left-hand side, $\mid Y_{k}\mid-\mid C_{s}^{k}(Y^{k-1},q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}))\mid$, is the difference between the number of contracts chosen when the capacity is (weakly) increased from $q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$ to $q^{k}_{s}(r_{1},...,r_{k-1})$. Hence, it cannot exceed the increase in capacity, which is $q^{k}_{s}(r_{1},...,r_{k-1})-q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$. Therefore, we now have
$\mid Y_{k}\mid-\mid Z_{k}\mid\leq q^{k}_{s}(r_{1},...,r_{k-1})-q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1}).$
Rearranging gives us
$q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})-\mid Z_{k}\mid\leq q^{k}_{s}(r_{1},...,r_{k-1})-\mid Y_{k}\mid.$
That is $\tilde{r}_{k}\leq r_{k}$.
We now prove (2). Consider a contract $x\in Y^{k}$. That is, $x\notin C_{s}^{k}(Y^{k-1},q^{k}_{s}(r_{1},...,r_{k-1}))$. When the set $Y^{k-1}$ is expanded to $\tilde{Y}^{k-1}$, $x$ cannot be chosen since the choice function satisfies the substitutes condition. That is,
$x\notin C_{s}^{k}(Z^{k-1},q^{k}_{s}(r_{1},...,r_{k-1})).$
When the capacity is reduced to $q^{k}_{s}(\tilde{r}_{1},...,\tilde{r}_{k-1})$, the quota monotonicity property implies that
$x\notin C_{s}^{k}(Z^{k-1},q^{k}_{s}(\widetilde{r}_{1},...,\widetilde{r}_{k-1})).$
That is, $x$ cannot be chosen in category $k$. Hence, $x\in Z^{k}$. Thus, we conclude that $Y^{k}\subseteq Z^{k}$.
Now, let $\tau_{j}=r_{j}-\tilde{r}_{j}$. As shown above, $\tau_{j}\geq 0$ for all $j=1,...,K_{s}$. Putting in $r_{j}=q_{s}^{j}(r_{1},...,r_{j-1})-\mid Y_{j}\mid$ and $\tilde{r}_{j}=q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{k-1})-\mid Z_{j}\mid$ gives
$\mid Z_{j}\mid=q_{s}^{j}(r_{1},...,r_{j-1})-q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})+\mid Y_{j}\mid+\tau_{j}.$
Summing both the right- and left-hand sides for $j=1,...,K_{s}$ yields
$\sum_{j=1}^{K_{s}}{\mid Z_{j}\mid}=\sum_{j=1}^{K_{s}}{\mid Y_{j}\mid}+\sum_{j=2}^{K_{s}}[q_{s}^{j}(r_{1},...,r_{j-1})-q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})]+\sum_{j=1}^{K_{s}}{\tau_{j}}.$
Since each $\tau_{j}\geq 0$, we have
$\sum_{j=1}^{K_{s}}{\mid Z_{j}\mid}\geq\sum_{j=1}^{K_{s}}{\mid Y_{j}\mid}+\sum_{j=2}^{K_{s}}{[q_{s}^{j}(r_{1},...,r_{j-1})-q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})]}.$
Since $q_{s}^{j}(r_{1},...,r_{j-1})\geq q_{s}^{j}(\tilde{r}_{1},...,\tilde{r}_{j-1})$ for all $j=2,...,K_{s}$, we obtain
$\sum_{[j=1}^{K_{s}}{\mid Z_{j}\mid}\geq\sum_{j=1}^{K_{s}}{\mid Y_{j}\mid}.$
That is, $\mid\overline{C}_{s}(Y;q_{s})\mid\leq\mid\overline{C}_{s}(Z;q_{s})\mid$, the desired conclusion.
Proof of Proposition 1
We need to show that $C_{s}^{mT}$ is in the GS family. The transfer functions under $C_{s}^{mT}$ satisfy monotonicity in a straightforward way. Fleiner (2001) shows that $C^{m}$ satisfies the substitute property Sönmez and Yenmez (2022) prove that $C^{m}$ satisfies size monotonicty.
We then only need to show that $C^{m}$ satisfies quota-monotonicity. That is, for any given set of applicants $A\subseteq\mathcal{I}$ and $q\in\mathbb{N}$, we need to show that
  1. (1)
    $C^{m}\left(A,q\right)\subseteq C^{m}\left(A,q+1\right)$, and
  2. (2)
    $\mid C^{m}\left(A,q+1\right)\mid-\mid C^{m}\left(A,q\right)\mid\leq 1$.
We consider two cases:
Case 1: The additional position is not reserved horizontally.
In this case, since the set of individuals $A\subseteq\mathcal{I}$ and the set of horizontally reserved positions $H^{v}=\bigcup_{j=1}^{L}H_{j}^{v}$ remain unchanged, when the capacity is increased, the horizontal type matching remains the same in step 1 of $C^{m}$. Moreover, since in Step 2 the remaining positions are filled following the merit scores ranking with individuals among the unassigned individuals until all positions are assigned or all individuals are selected, every individual who was chosen when capacity was $q$ will be selected when capacity increases to $q+1$. Therefore, both conditions of quota monotonicity are trivially satisfied.
Case 2: The additional position is reserved horizontally.
Without loss of generality, let $h_{j}$ be the horizontal type whose capacity increases by 1. For each $i\in C^{m}\left(A,q\right)$, given the applicants chosen before $i$ in the first step of $C^{m}$, $i$ is chosen because it increases the utilization of horizontally reserved positions, which means that horizontally reserved positions are available for her. Adding one more type $h_{j}$ position (weakly) increases the horizontally reserved positions available to her. Thus, $i$ continues to be chosen. That is, $C^{m}\left(A,q\right)\subseteq C^{m}\left(A,q+1\right)$. Since each applicant chosen when capacity was $q$ continues to be chosen when capacity increases by one, the number of applicants chosen in Step 1 of $C^{m}$ stays the same or increases by one. By Step 2 of $C^{m}$, we can conclude that $\mid C^{m}\left(A,q+1\right)\mid-\mid C^{m}\left(A,q\right)\mid\leq 1$.